A Complete Bibliography of Publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches

sur les pêcheries du Canada (1930–1939)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

Achromobacter [24]. Acid [32, 19]. Acids [44, 7]. Addition [37]. Adequate [2]. Adsorption [16]. Age [22]. Amino [44]. Among [17]. Analysis [49]. anguillaris [33]. Annual [3]. Appearance [27]. Approximate [7]. Artificial [14]. Atlantic [39, 34, 46, 11, 43, 31].

B.C. [5]. Bacteria [19]. Bacterial [38, 23]. Barren [13]. Behaviour [1]. Benzoic [32]. Bideford [26]. bifurcata [5]. Black [48]. Bodies [15]. Breakdown [44]. British [14, 3, 49, 13, 22].

c [7]. caerulea [7]. Canadian [25, 39, 11]. Canned [15]. Carbohydrates [44]. Clupea [47]. Coast [39]. Cod [8, 44, 42, 48]. Coefficients [12]. Columbia [14, 49, 13, 22, 3]. Composition [7]. Compound [37]. Concentration [19]. Condition [49, 47]. Conservation [36]. Content [21]. Cultus [14, 18, 3]. Current [50].

D [30, 41, 42, 48]. Descent [31]. Determining [47]. Development [10]. During [44, 17, 8]. Dynamic [15].

Eastern [25]. Effect [4]. Effectiveness [32]. Effects [50, 35]. Efficiencies [14]. Embryo [33]. Equilibrium [12]. Experiments [15].

Factor [49]. Factors [30, 31]. Fatness [47]. Fatty [7]. Features [45]. Feeding [40, 4]. Firmness [15]. Fish [6, 8, 21, 44, 7, 12, 36, 32, 23, 24]. Fluctuations [9]. Food [4, 29]. formed [16]. Frais [36]. French [36]. Fresh [36].

Gadus [42]. Girard [30]. gorbuscha [13, 22]. Gravity [47]. Gray [42]. Grayfish [30]. Gulf [34].

Halibut [41]. Harbour [5, 45]. Height [4]. Heritable [1]. Herring [49, 35, 40, 47]. Hippoglossus [41]. Histidine [37]. Homing [22]. Hydrogen [19]. Hydrogen-ion [19]. Hydrographic [45].

Ice [32]. II [8]. III [21, 12]. Influence [19]. Influencing [30, 31]. inlet [13]. Intensity [9]. Intestinal [41, 42, 48]. Investigation [14]. ion [19]. IV [23].

John [45]. Juice [8, 44].

Kelts [11]. Kingfishers [4]. Kokanee [18].

Ladysmith [5]. Lake [14, 17, 2, 3, 9, 18]. Larval [26, 10]. Length [46]. Life [26]. Light [35]. Ling [48]. Liver [30, 41, 42, 48]. Local [1].

macrocephalus [42]. Margaree [29]. Masset [13]. Maturity [22]. Measurement [15]. Mergus [29]. Method [15, 47]. Migrations [34]. Moisture [12]. molluscan [5]. Mortality [17]. Movement [50]. Movements [11, 35]. Multiple [15]. Multiple-Needle [15]. Muscle [8, 44, 12]. Muscles [21]. Mysidacea [25].

N.S. [29]. Natural [14]. Needle [15]. nerka [14, 17]. Nesting [33]. Nestlings [4]. Net [2, 3]. Nova [21].

OBITUARY [20, 28]. Oil [6, 7, 30]. Oils [6, 16, 7, 41, 42, 48]. Oncorhynchus [14, 17, 13, 22]. Order [27]. Origin [8]. Ostrea [26]. other [15]. Oxide [21, 38, 23, 24]. Oyster [26].

pallasii [47]. Pandalus [10]. Paphia [5]. Paul [9]. Pelagic [2, 3]. Penetrometer [15]. pH [36]. Pigments [6]. Pilchard [6]. Pilchards [7]. Pink [13, 22]. Plankton [2, 3]. Poisson [36]. Potencies [41, 42, 48]. Potency [30]. Preservation [32]. Press [8, 44]. Produced [8]. Propagation [14]. Proteins [44]. Proximate [49].

Quantitative [2]. Quantity [3].

Rôle [36]. Rainbow [9]. Red [48]. Reduction [38, 23]. Relation [49]. Relationship [46]. Relative [14]. Residence [17]. Residual [18]. Resistance [43]. Resistant [19]. Respiration [24]. Rigid [15]. River [29, 26]. Role [36, 24]. Runs [9].

Saint [45]. Salmon [15, 14, 17, 46, 11, 43, 13, 22, 1, 31, 18]. Salt [12]. Sampling [2]. Sardinops [7]. Scales [27]. Scotia [21]. Sea [11, 43]. Seaboard [34]. Season [49]. Seasonal [3]. Semi [15]. Semi-Rigid [15]. serrator [29]. situ [16]. Smolts [31]. Soaps [16]. Sockeye [14, 17]. Some [19]. Spawning [9]. species [5]. Specific [47]. Speckled [50, 27]. Spoilage [8, 21, 44, 23, 24]. Squalus [30]. Stages [17]. stenolepis [10]. Stream [34]. Studies [8, 21, 44, 12, 23, 24]. sucklii [30]. Surface [39].

Temperatures [39]. Temporary [34]. Tendency [22]. Transplantation [13]. Trend [17]. Trimethylamine [8, 21, 38, 23, 24]. Trout [50, 27, 9].

V [24]. Variations [3]. Various [17]. VI [44]. VII [6]. VIII [7]. virginica [26]. Vitamin [16, 30, 41, 42, 48].

Water [39, 34, 43, 4]. Waters [45, 25]. Weight [46].

Years [13]. Young [17].

Zoarces [33].

References

White:1938:LBS

[1] H. C. White and A. G. Huntsman. Is local behaviour in salmon heritable? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1):1–18, January 1938. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

    doi/full/10.1139/f38-001.

Ricker:1938:AQS

[2] William E. Ricker. On adequate quantitative sampling of the pelagic net plankton of a lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1): 19–32, January 1938. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f38-002.

Ricker:1938:SAV

[3] William E. Ricker. Seasonal and annual variations in quantity of pelagic net plankton, Cultus Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1):33–47, January 1938. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f38-003.

White:1938:FKF

[4] H. C. White. The feeding of kingfishers: Food of nestlings and effect of water height. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1):48–52, January 1938. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-004.

Quayle:1938:PBN

[5] Daniel B. Quayle. Paphia bifurcata, a new molluscan species from Ladysmith Harbour, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1): 53–54, January 1938. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f38-005.

Bailey:1938:FOV

[6] B. E. Bailey. Fish oils: VII. The pigments of pilchard oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1):55–58, January 1938. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-006.

Brocklesby:1938:FOV

[7] H. N. Brocklesby and K. F. Harding. Fish oils: VIII. The approximate composition of the fatty acids c the oil of pilchards (Sardinops caerulea). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(1):59–62, January 1938. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-007.

Beatty:1938:SFS

[8] S. A. Beatty. Studies of fish spoilage: II. The origin of trimethylamine produced during the spoilage of cod muscle press juice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):63–68, February 1938. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-008.

Mottley:1938:FIS

[9] C. McC. Mottley. Fluctuations in the intensity of the spawning runs of rainbow trout at Paul Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):69–87, February 1938. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-009.

Needler:1938:LDP

[10] Alfreda Berkeley Needler. The larval development of Pandalus stenolepis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):88–95, February 1938. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-010.

Huntsman:1938:SMC

[11] A. G. Huntsman. Sea movements of Canadian Atlantic salmon kelts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):96–135, February 1938. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-011.

Cooper:1938:SSF

[12] D. LeB. Cooper. Studies on salt fish: III. Equilibrium moisture coefficients of salt fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):136– 140, February 1938. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-012.

Pritchard:1938:TPS

[13] A. L. Pritchard. Transplantation of pink salmon (Oncorhynchus gorbuscha) into masset inlet, British Columbia, in the barren years. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(2):141–150, February 1938. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-013.

Foerster:1938:IRE

[14] R. E. Foerster. An investigation of the relative efficiencies of natural and artificial propagation of sockeye salmon (Oncorhynchus nerka) at Cultus Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3): 151–161, March 1938. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f38-014.

Charnley:1938:MFC

[15] F. Charnley and R. S. Bolton. The measurement of firmness of canned salmon and other semi-rigid bodies by the dynamic penetrometer method: I. Experiments with a multiple-needle penetrometer. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3):162–173, March 1938. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/f38-
015.

Brocklesby:1938:AVO

[16] H. N. Brocklesby and C. C. Kuchel. The adsorption of vitamin A from oils by soaps formed in situ. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3):174–183, March 1938. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-016.

Foerster:1938:MTA

[17] R. E. Foerster. Mortality trend among young sockeye salmon (Oncorhynchus nerka) during various stages of lake residence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3):184–191, March 1938. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-017.

Ricker:1938:RKS

[18] W. E. Ricker. “Residual” and Kokanee salmon in Cultus Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3):192–218, March 1938. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-018.

Watson:1938:IHI

[19] Dennis W. Watson. The influence of hydrogen-ion concentration on some acid resistant bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4a(3):

219–227, March 1938. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f38-019.

Anonymous:1939:Oa

[20] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4), April 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-020.

Beatty:1939:SFSa

[21] S. A. Beatty. Studies of fish spoilage: III. The trimethylamine oxide content of the muscles of Nova Scotia fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):229–232, April 1939. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-021.

Pritchard:1939:HTA

[22] A. L. Pritchard. Homing tendency and age at maturity of pink salmon (Oncorhynchus gorbuscha) in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):233–251, April 1939. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/f38-
022.

Watson:1939:SFSa

[23] Dennis W. Watson. Studies of fish spoilage: IV. The bacterial reduction of trimethylamine oxide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4): 252–266, April 1939. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f38-023.

Watson:1939:SFSb

[24] Dennis W. Watson. Studies of fish spoilage: V. The role of trimethylamine oxide in the respiration of achromobacter. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):267–280, April 1939. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f38-024.

Tattersall:1939:MEC

[25] W. M. Tattersall. The Mysidacea of Eastern Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):281–286, April 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-025.

Medcof:1939:LLO

[26] J. C. Medcof. Larval life of the oyster (Ostrea virginica) in Bideford River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):287–301, April 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-026.

Elson:1939:OAS

[27] Paul F. Elson. Order of appearance of scales in speckled trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(4):302–308, April 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-027.

Anonymous:1939:Ob

[28] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5), May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-028.

White:1939:FMS

[29] H. C. White. The food of Mergus serrator on the Margaree River, N.S. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):309–311, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-029.

Pugsley:1939:FIV

[30] L. I. Pugsley. Factors influencing the vitamin A and D potency of grayfish liver oil Squalus sucklii (Girard). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):312–322, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-030.

White:1939:FID

[31] H. C. White. Factors influencing descent of Atlantic salmon smolts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):323–326, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-031.

Tarr:1939:EBA

[32] H. L. A. Tarr and B. E. Bailey. Effectiveness of benzoic acid ice for fish preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):327– 336, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-032.

White:1939:NEZ

[33] H. C. White. The nesting and embryo of Zoarces anguillaris. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):337–338, May 1939. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-033.

Hachey:1939:TMG

[34] H. B. Hachey. Temporary migrations of Gulf Stream water on the Atlantic Seaboard. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):339– 348, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-034.

Johnson:1939:ELM

[35] W. H. Johnson. Effects of light on movements of herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):349–354, May 1939. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f38-035.

Nadeau:1939:PFR

[36] Aristide Nadeau. Le poisson frais: I. Le rôle du pH sur la conservation du poisson. (French) [Fresh fish: I. The role of pH on fish conservation]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):355–362, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-036.

Johnston:1939:ACH

[37] W. W. Johnston and M. L. Johnston. An addition compound of histidine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):363–366, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-037.

Tarr:1939:BRT

[38] H. L. A. Tarr. The bacterial reduction of trimethylamine oxide to trimethylamine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):367– 377, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-038.

Hachey:1939:SWT

[39] H. B. Hachey. Surface water temperatures of the Canadian Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):378–391, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-039.

Johnson:1939:FH

[40] W. H. Johnson. Feeding of the herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):392–395, May 1939. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f38-040.

Pugsley:1939:VDPa

[41] L. I. Pugsley. Vitamin A and D potencies of liver and intestinal oils of halibut (Hippoglossus hippoglossus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):396–404, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-041.

Pugsley:1939:VDPb

[42] L. I. Pugsley. Vitamin A and D potencies of liver and intestinal oils of Gray cod (Gadus macrocephalus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):405–408, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-042.

Huntsman:1939:RAS

[43] A. G. Huntsman and W. S. Hoar. Resistance of Atlantic salmon to sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):409– 411, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-043.

Beatty:1939:SFSb

[44] S. A. Beatty and V. K. Collins. Studies of fish spoilage: VI. The breakdown of carbohydrates, proteins and amino acids during spoilage of cod muscle press juice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5): 412–423, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-044.

Hachey:1939:HFW

[45] H. B. Hachey. Hydrographic features of the waters of Saint John Harbour. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):424–440, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-045.

Hoar:1939:WLR

[46] William S. Hoar. The weight–length relationship of the Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):441–460, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-046.

Tester:1939:SGM

[47] Albert L. Tester. A specific gravity method for determining fatness (condition) in herring (Clupea pallasii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):461–471, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f38-047.

Pugsley:1939:VDPc

[48] L. I. Pugsley. Vitamin A and D potencies of liver and intestinal oils of red, black and ling cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5): 472–477, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-048.

Hart:1939:PAB

[49] John Lawson Hart, Albert L. Tester, Desmond Beall, and John P. Tully. Proximate analysis of British Columbia herring in relation to season and condition factor. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):478– 490, May 1939. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f38-049.

Elson:1939:ECM

[50] Paul F. Elson. Effects of current on the movement of speckled trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 4b(5):491–499, May 1939. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f38-050.